OpenStudy (anonymous):

What is the value of a? 9a - 3(a-6) = -6

OpenStudy (anonymous):

@sumanta

OpenStudy (anonymous):

9a - 3a+18 = -6

OpenStudy (anonymous):

Did u mean -18??

OpenStudy (anonymous):

so 6a = -18-6

OpenStudy (anonymous):

no

OpenStudy (anonymous):

9 x -18 + 3 = -6

OpenStudy (anonymous):

so 6a = -24

OpenStudy (anonymous):

yes a = - 4

OpenStudy (anonymous):

The length of a rectangle is 12 in. more than its width. The perimeter is 64 in. What is the length of the rectangle? in.

OpenStudy (anonymous):

so let width be x

OpenStudy (anonymous):

(x+12)* x = 64

OpenStudy (anonymous):

i dont even know

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

I am confused here

OpenStudy (anonymous):

22

OpenStudy (anonymous):

may be

OpenStudy (anonymous):

i got it right for 22 lol

OpenStudy (anonymous):

The easy way: deduct 12 (the difference) from 32 (the sum of length and width ie half the perimeter) and divide by 2 to find width. So width is 10 and length is 22.

OpenStudy (anonymous):

go for it